a pharmacist wants to mix an ointment that is 6% zinc with and ointment that is 12% zinc oxide to make 30 grams of an ointment that that is 10% zinc oxide. How many grams of each ointment should the pharmacist mix together ?
Answer by richwmiller(17219)

x+y=30,
0.06*x+0.12*y=0.1*30
0.06*x+0.12*y=3
x=30-y
0.06*(30-y)+0.12*y=3
1.8-0.06y+0.12*y=3
0.06*y=1.2
y=20 at 12%
x=30-y
x=10 at 6%
check
0.06*10+0.12*20=0.1*30
0.6+2.4=3
3=3
ok